OpenStudy (anonymous):

Distributive Property (show work) 1) 3(x + 2) = -9 -2x 2) 3(x + 2) = 9 +2 (x + 4) 3) 5(c + 4) = 4(2c - 3) -7

OpenStudy (bahrom7893):

3x + 6 = -9 - 2x 3x + 2x = -9-6 5x = -15 x = -3

OpenStudy (bahrom7893):

3x + 6 = 9 + 2x + 8 3x - 2x = 17-6 x = 11 (if i didn't make any dumb mistakes)

OpenStudy (bahrom7893):

no i didn't

OpenStudy (bahrom7893):

3) 5c + 20 = 8c - 12 - 7 5c - 8c = -20 - 19 -3c = -39 c = 13